Annals of the University of Craiova, Mathematics and Computer Science Series Volume 37(1), 2010, Pages 1–3 ISSN: 1223-6934

Instructions to Authors: This is the Title

First Author, Second Anonymous, and Third Fictive-Author

Abstract. The abstract appears at the beginning of the paper, and not exceeding 300 words. This paper contains brief instructions for authors who wish to prepare their manuscripts for AMI journal by means of LaTeX. You will find the format you have to choose, how to type the title of your paper, the titles of sections, examples of definitions, lemmas, theorems, equations etc. Submissions to Annals of the University of Craiova - Mathematics and Computer Science Series (AMI) must represent original material. Papers are accepted for review with the understanding that the same work has been neither submitted to, nor published in, another journal or conference.

2010 Mathematics Subject Classification. Primary 60J05; Secondary 60J20.

Key words and phrases. rules, figures, citation of papers, references, examples. (4–6 keywords.)

1. Introduction

Manuscripts should be written in English or French, following the style of our journal in what concerns the technical preparation of the papers. The manuscripts will include the full address (es) of the author(s), with E-mail address(es), an abstract not exceeding 300 words, 2010 Mathematics Subject Classification, Key words and phrases, and should be limited to 15,000 words. The name, affiliation (institution and department), regular mailing address and email of the author(s) should be filled in at the end of the paper. The manuscripts must be prepared electronically in LATEX macro package.

If it is determined that a paper has already appeared in anything more than a conference proceeding, or appears in or will appear in any other publication before the editorial process at AMI journal is completed, the paper will be automatically rejected. Concurrent submission to AMI journal and other publications is viewed as a serious breach of ethics and, if detected, will result in immediate rejection of the submission.

To ensure that all articles published in the journal have a uniform appearance, authors must produce a PDF document that meets the formatting specifications outlined here. The PDF document will be used for both the hardcopy and electronic versions of the journal. Papers must be printed in the single column format as shown in the enclosed sample.

Received January 11, 2010.

The presented work has been conducted in the context of the GRANT-xy funded by The National University Research Council from Romania, under the contract 01/2008.

2. Topics

The Annals of the University of Craiova - Mathematics and Computer Science series - publishes original research papers in any branch of Mathematics and Computer Science. The publishing policy of AMI journal encourages particularly the publishing of scientific papers that are focused on the convergence of the mathematics and computer science.

3. Definitions and Examples

This is the section named Definitions and Examples.

This is the equation 1.

x^{3} + y^{3} = z^{3}  (1)

This equation is not numbered.

x^{2} + y^{2} = z^{2}

Definition 3.1. This is the Definition 3.1.

Example 3.1. This is the Example 3.1.

Theorem 3.1. This is the Theorem 3.1.

Proof. This is the proof of Theorem 3.1. □

Lemma 3.1. This is Lemma 3.1.

Proof. This is the proof of Lemma 3.1. □

Proposition 3.1. This is the Proposition 3.1.

Proof. This is the proof of Proposition 3.1. □

Corollary 3.1. This is the Corollary 3.1.

Proof. This is the proof of Corollary 3.1. □

Remark 3.1. This is the Remark 3.1.

3.1. Citations. Citations within the text should be as shown in this document, for example ([1], [3], [4], [5], [6], [7]).

References to figures, tables, sections, examples, theorems and so on should be capitalized, as in “In Defnition 3.1 we show that ...”

4. Figures and Tables

Figures and tables should be inserted in proper places throughout the text. Any graphics or pictures must be saved in Encapsulated PostScript(.eps) format.

Latex handles figures that are EPS. EPS files are vector-scaled meaning that resizing will not decrease image quality - an essential aspect in camera-quality documents(see Fig 1). Examine this template and you’ll see that a EPS image has been included (a picture of an assembly work cell). The latex statements are:

Figure 1. Text is written below the image

4.1. Figure properties. The caption command assigns what text is written below the image fig.eps. The label command is used so that you can reference it within the paper. You’ll notice fig.eps is repeated twice. This was done just to show two pictures can be placed side-by-side. The width variable scales the figures’ height automatically.

5. Examples of the AMI format for References

Book

S. Russell and P. Norvig, Artificial Intelligence: A Modern Approach (Third Edition), Prentice Hall, 2009.

Book chapter

B. Brown and M. Aaron, The politics of mathematics. In J. Smith (ed) The rise of modern science, 3rd edition, Wiley, New York (2001), 230–257.

Journal Article

O. Goubet, Ondes hydrodynamiques amorties, Annals of the University of Craiova, Mathematics and Computer Science Series 32 (2005), no. 1, 40–51.

Proceedings Paper

C.I. Popescu, P.-C. Luis and H. Alonso, A Java Implementation of Mobile Agents, Proceedings of the International Conference on Multi-Agents Systems, Barcelona, Spain, June 14-17, 2004 (Garcia-Falset, J. et al., Eds.), IEEE Press, Los Alamitos (2005), 31-38.

or

P. Thati, P.-H. Chang and G. Agha, Crawlets: Agents for high performance web search engine, International Conference on Mobile Agents(MA 2005), Craiova, Romania, July 21-23, 2005, Lecture Notes in Computer Science 2240, Springer-Verlag, Berlin (2006), 19-34.

Univeristy Technical Report

I. Mierswa, Making indefinite kernel learning practical, Technical Report TR 41-06, University of Dortmund (2006).

Dissertation or Thesis

W.J. Clancey, Transfer of rule-based expertise through a tutorial dialogue, Ph.D. thesis, Department of Computer Science, Stanford University (1979).

Online document/URL Reference

  http://en.wikipedia.org/wiki/Remote_procedure_call

6. Conclusion

Please be aware that editors and reviewers are required to check the submitted manuscript to determine whether a sufficient amount of new material has been added to warrant publication in AMI journal. If you have used your own previously published material as a basis for a new submission, then you are required to cite the previous work(s) and clearly indicate how the new submission offers substantively novel or different contributions beyond those of the previously published work(s). Any manuscript not meeting these criteria will be rejected.

Papers must be submitted online. You should send us the LaTex source file and the graphics. You must send us also the pdf version of your paper. Manuscripts must be accompanied by a signed copyright transfer form. The copyright transfer form is

available for download at http://inf.ucv.ro/~ami

The reference section should be labeled “References ”and should appear at the end of the paper. A sample list of references is given at the end of these instructions. Poorly prepared, incomplete or sloppy references reflect badly on the quality of your paperwork and your research. Please prepare complete and accurate citations.

References should be listed in alphabetical order; the following reference style should be used:

References

[1] S. Russell and P. Norvig, Artificial Intelligence: A Modern Approach (Third Edition), Prentice Hall, 2009.

[2] B. Brown and M. Aaron, The politics of mathematics. In J. Smith (ed) The rise of modern science, 3rd edition, Wiley, New York (2001), 230–257.

[3] O. Goubet, Ondes hydrodynamiques amorties, Annals of the University of Craiova, Mathematics and Computer Science Series 32 (2005), no. 1, 40–51.

[4] C.I. Popescu, P.-C. Luis and H. Alonso, A Java Implementation of Mobile Agents, Proceedings of the International Conference on Multi-Agents Systems, Barcelona, Spain, June 14-17, 2004 (Garcia-Falset, J. et al., Eds.), IEEE Press, Los Alamitos (2005), 31-38.

[5] I. Mierswa, Making indefinite kernel learning practical, Technical Report TR 41-06, University of Dortmund (2006).

[6] W.J. Clancey, Transfer of rule-based expertise through a tutorial dialogue, Ph.D. thesis, Department of Computer Science, Stanford University (1979).

[7] http://en.wikipedia.org/wiki/Remote_procedure_call

(First Author, Second Anonymous) Department of Computer Science, University of Craiova, 13 A.I. Cuza Street, Craiova, 200585, Romania

E-mail address: first.author@inf.ucv.ro, second nonymous@yahoo.com

(Third Author) Department of Mathematics, University of Craiova, 13 A.I. Cuza Street, Craiova, 200585, Romania

E-mail address: third.fictive@inf.ucv.ro